# PIGEON: Predicting Image Geolocations

### 17 Dec 2023 v4

@@ Appendix F Deployment to Geoguessr
F.1 Game mode
Geoguessr can be played in both single and multi-player modes. In our performance evaluation of PIGEON, we decided to focus on Geoguessr’s Competitive Duels mode, whereby the user directly competes with an opponent in a multi-round game with increasing round difficulty. Each guess is translated into a Geoguessr score whose formula we reverse-engineered by recording results from the game. The formula for the Geoguessr score on the world map is approximately
@@ F.2 Chrome extension
    Optional: Random delay to behave more human-like.
@@ F.3 Inference API
  • Inference endpoint. A POST endpoint that receives either one or four images, passes them through a preprocessing pipeline and then runs inference on a GPU. In addition, it saves the images temporarily on disk for later evaluation. Finally, the API returns the latitude & longitude predictions of PIGEON to the client.
  • Statistics endpoint. A POST endpoint that receives the statistics about the correct location, the score & distance of our guess, and human performance. This data is saved on disk and later used in our evaluations.
Consequently, our work demonstrates that PIGEON can effectively be applied in real-time scenarios as a system capable of end-to-end planet-scale image geolocalization.
sixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 07 Apr 2024 v5

@@ Appendix H Deployment to GeoGuessr
H.1 Data overlap in live games
When deploying PIGEON in live games against human players, controlling for locations not in PIGEON’s training dataset is impossible. Across all live games from Figure 4, we find that $1.8\%$ of game locations are within less than $100$ meters of any location in our training dataset. This slight overlap between training and live evaluation locations is not problematic because top human players would see more unique locations over the course of their GeoGuessr career, resulting in an even larger overlap between already seen and new live data for them.
H.2 Game mode
GeoGuessr can be played in both single and multi-player modes. In our live performance evaluation of PIGEON, we decided to focus on GeoGuessr’s Competitive Duels mode, whereby the user directly competes with an opponent in a multi-round game with increasing round difficulty. Notably, while our GeoGuessr bot simply takes four images spanning the entire GeoGuessr panorama, other players can additionally move around in the Street View scene for at least 15 seconds which is the minimum time available to the opponent once a guess is made, resulting in them gathering more relevant information to refine their prediction. Each guess is subsequently translated into a GeoGuessr score whose formula we reverse-engineered by recording results from the game. The formula for the GeoGuessr score on the world map is approximately:
@@ H.3 Chrome extension
    Optional: Random delay before making a guess to make the model’s behaviour seem more human-like.
@@ H.4 Inference API
  • Inference endpoint. A POST endpoint that receives either one or four images, passes them through a preprocessing pipeline and then runs inference on a GPU. In addition, it saves the images temporarily on disk for later evaluation. Finally, the API returns the latitude and longitude predictions of PIGEON to the client.
  • Statistics endpoint. A POST endpoint that receives the statistics about the correct location, the score and distance of our guess, and human performance. This data is saved on disk and later used to generate summary statistics.
Our work demonstrates that PIGEON can effectively be applied in real-time scenarios as a system capable of end-to-end planet-scale image geolocalization.
Appendix I Acknowledgements
We extend our gratitude to Erland Ranvinge, the Chief Technical Officer of GeoGuessr. Mr. Ranvinge’s expertise and willingness to share in-depth knowledge about the GeoGuessr web API was instrumental in deploying PIGEON to GeoGuessr. His support in facilitating live evaluations against human players not only enhanced the quality of our evaluations but also provided unique insights that enriched our findings. We further want to highlight that this work was initially conceived in an independent class project as part of CS 330: Deep Multi-Task and Meta Learning, taught by Professor Finn at Stanford University, and subsequently evolved into the comprehensive study presented in this work. Finally, we wish to express our appreciation to Stanford’s faculty and administrators who supported our work through cloud and compute resources.
Figure 13: Example predictions of PIGEOTTO on fourteen images from the YFCC26k Müller-Budack et al. (2018) test set.